OpenStudy (anonymous):

NEED HELP!!!!! Which of these sequences is a geometric sequence? A. 1, 2, 4, 7, 11, 16, 22, … B. 2, 4, 8, 14, 22, 38, … C. 3, 6, 9, 12, 15, 18, 21, … D. 3, 9, 27, 81, 343, 729, …

Mehek (mehek14):

which one do you think?

OpenStudy (anonymous):

c

Mehek (mehek14):

it is not C because they are not being multiplied by the same number \(3*2 = 6\\ 6*1.5 = 9\) and so on

Mehek (mehek14):

a geometric sequence is where each term after the first is found by multiplying the previous one by the same number

OpenStudy (anonymous):

A! RIGHT

Mehek (mehek14):

\(1*2=2\\2*2=4\\4*2=8~not~7\)

Mehek (mehek14):

the pattern in A is \(1+1=2\\2+2=4\\4+3=7\\7+4=11\)

OpenStudy (anonymous):

d!! new I should of whent with my gut

Mehek (mehek14):

yes D they increase by \(*3\) each time

OpenStudy (anonymous):

thx I have more questions though can you help me with them

Mehek (mehek14):

sure post separately

OpenStudy (anonymous):

k